# Residual Rotation Correction using Tactile Equivariance

### v1

@@ III-B Flow Matching Policy
At test time, the policy generates actions by sampling $x_{0}\!\sim\!\mathcal{N}(0,\sigma^{2}I)$ and numerically integrating the learned ODE $\frac{dx}{dt}=v_{\theta}(x,t,c)$ with initial condition $x(0)=x_{0}$. The integration is performed from $t=0$ to $t=1$ using NN Euler steps: $x_{i+1}=x_{i}+\frac{1}{N}\,v_{\theta}(x_{i},t_{i},c)$ where $t_{i}=\frac{i}{N}$, yielding the final action $x_{N}$.
@@ IV-D Equivariant Angular Residual Tracking
After acquiring the normal map NN, we aim to train an equivariant network $\phi:N\mapsto(\cos-\alpha_{t},\sin-\alpha_{t})$ that predicts the target in-hand yaw $\alpha_{t}$ in the form of a unit vector on a unit circle. The model satisfies
@@
where $R_{g}$ is the standard $2\times 2$ rotation matrix. Thus, when the object rotates on the finger by $g\in\mathrm{SO}(2)$, the input normal map NN becomes $g\cdot N$, and the output target in-hand yaw vector will counter-rotate accordingly, compensating for the in-hand object rotation. This property facilitates zero-shot generalization, achieving precise angular estimation without training on tactile images with different object orientations.

### v2

@@ III-B Flow Matching Policy
At test time, the policy generates actions by sampling $x_{0}\!\sim\!\mathcal{N}(0,\sigma^{2}I)$ and numerically integrating the learned ODE $\frac{dx}{dt}=v_{\theta}(x,t,c)$ with initial condition $x(0)=x_{0}$. The integration is performed from $t=0$ to $t=1$ using $N$ Euler steps: $x_{i+1}=x_{i}+\frac{1}{N}\,v_{\theta}(x_{i},t_{i},c)$ where $t_{i}=\frac{i}{N}$, yielding the final action $x_{N}$.
@@ IV-D Equivariant Angular Residual Tracking
After acquiring the normal map $N$, we aim to train an equivariant network $\phi:N\mapsto(\cos-\alpha_{t},\sin-\alpha_{t})$ that predicts the target in-hand yaw $\alpha_{t}$ in the form of a unit vector on a unit circle. The model satisfies
@@
where $R_{g}$ is the standard $2\times 2$ rotation matrix. Thus, when the object rotates on the finger by $g\in\mathrm{SO}(2)$, the input normal map $N$ becomes $g\cdot N$, and the output target in-hand yaw vector will counter-rotate accordingly, compensating for the in-hand object rotation. This property facilitates zero-shot generalization, achieving precise angular estimation without training on tactile images with different object orientations.